Suppose that you are in Singapore with 10,000 Singapore dollars (SGD) and you want to exchange them for euros for your trip to Paris. The exchange shop gives you the following rates: Commission: 1% USD/SGD = 1.2375 EUR/USD = 1.2800. How many dollars will you get for your Singapore dollars?
Answer #1

You want to exchange 10,000 Singapore dollars (SGD) to Euros and

Exchange rates are given

USD/SGD = 1.2375. This means that 1 USD = 1.2375 SGD

Therefore, 1 SGD = (1/1.2375) USD = 0.808081

Now, the exchange shop charges 1% as commission. So commission charged = 1% * 10000 SGD = 100 SGD

So, you are left with 10000 - 100 = 9900 SGD

9900 SGD = 9900 * 0.808081 USD = 8000 USD

Hence we will get 8000 USD for 9900 SGD

Answer -> 8000 USD